# Scaling laws for velocity profile of granular flow in rotating drums

### v2

@@ Abstract
We numerically analyze the steady flow of two-dimensional granular materials in a rotating drum using the discrete element method and a continuum model. The velocity fields obtained from both methods are quantitatively consistent. Two distinct regions exist in the granular flow: the surface flow layer and the static flow regime associated with the rigid rotation near the bottom. The thickness of the surface flow layer increases with the drum diameter and shows a slight dependence on the angular velocity of the rotating drum. We derived scaling laws for the velocity profile and surface layer thickness using dimensional analysis applied to the continuum equations. The validity of the scaling laws is confirmed numerically.
@@ I Introduction
Flows in rotating drums are used in various industrial processes of granular materials, including mixing, breaking, granulating, drying, and chemical reactions. The form of the flow inside the drum depends on the Froude number $Fr=\Omega^{2}D/2g$ [1], where $\Omega$ is the angular velocity of the rotation, $D$ is the diameter of the drum, and $g$ is the gravitational acceleration. For $10^{-4}<Fr<10^{-2}$, a steady flow driven by the rotating drum occurs, as shown in Fig. 1. In this steady state, the flow field comprises two regions: the “surface flow layer” and the “static flow regime.” In the surface flow layer, the granular material flows down along the slope of the free surface. In the static flow regime, the granular material exhibits a solid-like rigid rotation following the drum wall.
Numerous experimental studies have investigated the dependence of the surface flow layer thickness $h$ on $D$ and $\Omega$ in drums half-filled with granular materials [2, 3, 4, 5, 6]. In. Ref. [2], it is reported that $h$ is proportional to $D$ and almost independent of $\Omega$ in a drum long enough in the axial direction. According to Ref. [5], $h$ in axially thin rotating drums exhibits a power-law dependence on $\Omega$ for sufficiently large $D/d$ with the particle diameter $d$. The experiments in Ref. [7] suggested that $h$ in thin drums obeys $h/d\propto\tilde{Q}^{0.44}$ with the dimensionless flow rate $\tilde{Q}=\frac{1}{2}Fr^{1/2}\left(D/d\right)^{3/2}$. However, the dependence of $h$ has not been explained theoretically.
@@
[…]
Figure 1: Schematic of the steady granular flow in a rotating drum.
Numerical simulations have been conducted to analyze the granular flow in rotating drums. Several researchers have used the discrete element method (DEM) [8, 9, 10, 11, 12], where the equation of motion for each particle is solved numerically. However, the numerical cost for the DEM simulation increases with the number of particles in the system with $D\gg d$. Therefore, direct simulations of the systems used in industrial applications with a large number of particles are challenging. In recent years, simulation methods based on continuum equations have been proposed, which are free from the restriction on the number of particles. In Refs. [13, 14, 15], granular materials are treated as elastoplastic bodies to investigate their mixing driven by rotating drums. In Refs. [16, 17], the computational fluid dynamics (CFD) method was used to study the flows in rotating drums, where granular materials are treated as non-Newtonian viscous fluids. It has been reported that the velocity field in the rotating drums is reproduced by these continuum approaches; however, its dependence on $D$ and $\Omega$ has not been studied in detail.
Various constitutive relations have been utilized in these continuum approaches. Among the constitutive relations, the $\mu(I)$-rheology has been applied to various granular flows because of its simplicity [18, 19]. In the $\mu(I)$-rheology, the shear stress $\tau$ of granular materials under uniform shear with the shear rate $\dot{\gamma}$ and pressure $p$ is given by
$\tau=\mu(I)p.$ (1)
Here, $\mu(I)$ is the bulk friction depending on the inertial number $I=\dot{\gamma}/\sqrt{p/\rho_{s}d^{2}}$, where $\rho_{s}$ is the density of the particles comprising the granular materials. In Ref. [19], a functional form of $\mu(I)$ is proposed as follows:
$\mu\left(I\right)=\mu_{s}+\frac{\mu_{2}-\mu_{s}}{I_{0}/I+1},$ (2)
where $\mu_{s}$, $\mu_{2}$, and $I_{0}$ are constants depending on the elasticity, dissipation, and friction of the particles [18]. The continuum model using the $\mu(I)$-rheology treats granular materials as non-Newtonian fluids with the viscosity $\eta=\mu(I)p/\dot{\gamma}$. Recently, the constitutive relations incorporating non-local effects in the $\mu(I)$-rheology have been proposed [20, 21, 22]. The continuum models using the $\mu(I)$-rheology reproduce various granular flows, including those in rotating drums [23, 13, 14, 16]. Therefore, this constitutive law is expected to explain the dependence of the velocity profile on $D$ and $\Omega$.
This study aims to understand the steady flow of granular materials in rotating drums. We focus on 2D rotating drums half-filled with granular materials for $D/d\gg 1$ and obtain the velocity field numerically by DEM and CFD simulations. The details of the numerical simulations are described in Sec. II. The dependence of the velocity field on $D$ and $\Omega$ is shown in Sec. III, where we demonstrate that the continuum model quantitatively reproduces the flow obtained in the DEM simulation over a broad parameter space. In Sec. IV, scaling laws characterizing the velocity field are derived analytically from the continuum model using the $\mu(I)$-rheology and verified numerically. The method for calculating the velocity field in the DEM simulation is presented in Appendix A. The values of the parameters $\mu_{s}$, $\mu_{2}$, and $I_{0}$ for the $\mu(I)$-rheology are estimated in Appendix B. The packing fraction of granular materials in rotating drums is demonstrated in Appendix C. The position of the free surface is discussed in Appendix D. The details of the derivation for the analytical expression of $h$ are shown in Appendix E.
II Numerical Method
II.1 Setup
In this paper, we consider a two-dimensional granular flow in a rotating drum, as illustrated in Fig. 1. The origin of the Cartesian coordinate system $(X,Z)$ is placed at the center of the drum with diameter $D$. The polydisperse particles with the minimum diameter $d$ are half-filled in the drum. The drum is rotated counterclockwise with the angular velocity $\Omega$. The drum diameter $D$ and angular velocity $\Omega$ are set to $D=150d$, $300d$, and $450d$ and $1.26\times 10^{-3}\leq\Omega/\sqrt{g/d}\leq 1.26\times 10^{-2}$ to satisfy $10^{-4}<Fr<10^{-2}$, respectively.
II.2 DEM simulation
II.2.1 Time evolution equations

### v3

@@ Abstract
We theoretically and numerically investigate the steady flow of two-dimensional granular materials in a rotating drum using the discrete element method and a continuum model with the $\mu(I)$-rheology. The velocity fields obtained from both methods are in quantitative agreement. The granular flow exhibits two distinct regions: a surface flow layer and a static flow regime corresponding to rigid rotation near the drum bottom. The thickness of the surface flow layer increases with the drum diameter and shows a weak dependence on the angular velocity of the drum. Using dimensional analysis of the continuum equations, we analytically identify nondimensional parameters for the velocity profile and the surface flow layer thickness, which lead to scaling laws characterising the flow in rotating drums with low Froude number and large system size. The validity of the scaling laws is confirmed by numerical simulations.
@@ I Introduction
Granular flows in rotating drums play a central role in a wide range of industrial processes, such as mixing, size reduction, granulation, drying, and chemical reactions. The flow regime inside the drum is primarily characterized by the Froude number, $Fr=\Omega^{2}D/2g$, where $\Omega$ is the angular velocity, $D$ is the drum diameter, and $g$ is the gravitational acceleration. [1] For the rolling state with $10^{-4}<Fr<10^{-2}$, the system exhibits a steady flow [1], which can be divided into two distinct regions. Near the free surface, a thin “surface flow layer” is formed where particles continuously cascade down the slope. Beneath this layer, particles undergo a solid-body rotation with the drum wall, often referred to as the “static flow regime” (see Fig. 1).
Figure 1: Schematic of steady granular flow in a rotating drum. $X$ and $Z$ denote the horizontal and vertical coordinates, respectively, with the origin located at the drum center. The $x$-axis is taken parallel to the free surface, and the $z$-axis is normal to it.
Numerous experimental studies have examined how the thickness of the surface flow layer, $h$, in the rolling state depends on the drum diameter $D$ and angular velocity $\Omega$ in half-filled drums [2, 3, 4, 5, 6, 7]. For example, Parker et al. [2] studied axially long drums with glass beads and reported that $h$ scales linearly with $D$ and is nearly independent of $\Omega$. In contrast, Félix et al. [5] investigated thin drums and found that $h$ measured near the walls exhibits a power-law dependence on $\Omega$ when $D\gg d$. Pignatel et al. [8] further suggested that $h/d$ for glass and steel spheres in thin drums sandwiched between acrylic walls follows $h/d\propto\tilde{Q}^{0.44}$, where $\tilde{Q}=\tfrac{1}{2}Fr^{1/2}(D/d)^{3/2}$ is the dimensionless flow rate. Taken together, these studies indicate that the scaling of $h$ depends on the system geometry, sidewall effects, and material properties, and varies significantly across different Froude and inertial number regimes. However, a consistent theoretical explanation for these dependencies has not yet been established.
Granular flows in rotating drums have also been studied extensively using discrete element method (DEM) simulations, in which the equations of motion for individual particles are solved numerically. [9, 10, 11, 12, 13] These simulations have successfully reproduced the velocity profiles observed in experiments. However, direct simulations of systems at industrially relevant scales remain impractical, since the computational cost of DEM grows rapidly with the particle number when $D\gg d$. Moreover, while DEM provides detailed particle-level information, it does not by itself yield a theoretical description of the dependence of the surface flow layer thickness $h$ on $D$ and $\Omega$.
Another approach is to use continuum equations to investigate granular flows in rotating drums, providing a macroscopic description of the flow behavior. Among various constitutive relations proposed in this framework, the so-called $\mu(I)$-rheology has become widely used due to its simplicity and ability to capture dense granular flows. [14, 15, 16] In the $\mu(I)$-rheology, granular media is treated as a non-Newtonian fluid whose viscosity depends on the local shear rate and pressure, and extensions incorporating non-local effects have also been developed. [17, 18, 19] Continuum models based on such constitutive relations have successfully reproduced velocity profiles in rotating drums across several setups. [20, 21, 22, 23, 24, 25] Nevertheless, systematic studies on how these models capture the dependence of the surface flow layer thickness $h$ on $D$ and $\Omega$ are still lacking. In addition, most prior work has relied primarily on computational fluid dynamics (CFD) simulations of the continuum model to compute the velocity field numerically, rather than on analytical understanding.
For conventional fluids, theoretical understanding has been achieved by analyzing the continuum equations, i.e., the Navier-Stokes equations. [26] If continuum models for granular materials can successfully reproduce experimental and DEM velocity profiles, one may expect that similar theoretical analyses would clarify the dependence of the velocity field on system parameters such as $D$ and $\Omega$ in rotating drums. However, to the best of our knowledge, such a theoretical explanation has not yet been developed, despite the numerical evidence of the validity of continuum descriptions for granular flows.
This study aims to provide a theoretical understanding of steady granular flows in rotating drums. As a first step, we focus on two-dimensional drums half-filled with granular materials in the limit $D/d\gg 1$. This setting enables an analytically tractable formulation of the continuum model and remains directly relevant to experiments in sufficiently long drums, where the flow can be approximated as two-dimensional. We first demonstrate, by comparing with discrete element method (DEM) simulations, that the continuum model with the $\mu(I)$-rheology quantitatively reproduces the velocity field. Building on this validation, we then derive scaling laws for the velocity profile and the surface flow layer thickness $h$ from the continuum model and verify them against DEM results.
This paper is organized as follows. Section II describes the setup of our models. Section III presents the velocity fields obtained from DEM and continuum simulations. Scaling laws derived from the continuum model are given in Sec. IV. Sec. V summarizes the main findings and discusses possible directions for future work. Additional details, including the determination of rheological parameters and analytical derivations, are provided in the Appendices.
II Setup
In this study, we consider a two-dimensional granular flow in a rotating drum, as illustrated in Fig. 1. This choice captures the essential features of the flow relevant to testing the theoretical predictions, while keeping the computational cost manageable. The origin of the Cartesian coordinate system $(X,Z)$ is placed at the center of the drum with diameter $D$. Polydisperse particles with minimum diameter $d$ are half-filled in the drum. The drum rotates counterclockwise with angular velocity $\Omega$. We consider three drum diameters: $D=150d$, $300d$, and $450d$. The angular velocity is varied within $1.26\times 10^{-3}\leq\Omega/\sqrt{g/d}\leq 1.26\times 10^{-2}$ to satisfy $10^{-4}<Fr<10^{-2}$. To provide reference data for the continuum analysis, we first perform discrete element method (DEM) simulations in this 2D system. The particle-scale information obtained from DEM serves as a benchmark for validating the continuum model, which will be introduced in Sec. II.2 and serves as the primary tool for deriving theoretical insights into the velocity field.
II.1 DEM simulation
In the DEM simulation, [27, 28, 29] we consider $N$ particles of identical density $\rho_{s}$ in a drum. Particle interactions are modeled as linear springs, characterized by the normal elastic constant $k^{(\mathrm{n})}$, tangential elastic constant $k^{(\mathrm{t})}$, restitution coefficient $e$, and particle friction coefficient $\mu_{\mathrm{p}}$. The rough wall surface of the drum is modeled by placing particles of diameter $d$.
We use polydisperse particles, [11] with diameters uniformly distributed from $d$ to $1.2d$. The parameters are set as follows: $k^{(\mathrm{n})}=2.0\times 10^{5}mg/d$, $k^{(\mathrm{t})}/k^{(\mathrm{n})}=2/7$, $e=0.92$, and $\mu_{p}=0.50$, [30] where $m$ is the mass of a particle of diameter $d$. The time evolution equations are solved numerically using the leapfrog algorithm with time step $\Delta t=1.0\times 10^{-4}\sqrt{d/g}$. The drum is half-filled with $N$ particles, set to $7700$, $30600$, and $69200$ for $D=150d$, $300d$, and $450d$, respectively. Details of the DEM simulation are given in Appendix A. From the DEM simulation, we calculate the velocity field in the steady state using the procedure described in Appendix B.
II.2 Continuum model
To gain a theoretical understanding of the flow behavior observed in the DEM simulations, we adopt a continuum model that treats the granular bed in the rotating drum as an incompressible, non-Newtonian fluid. The governing equations consist of the mass and momentum conservation equations. The mass conservation equation is given by
$\nabla\cdot{\bf\it v}=0,$ (1)
and the momentum conservation equation reads
$\frac{\partial{\bf\it v}}{\partial t}=-\left({\bf\it v}\cdot\nabla\right){\bf\it v}-\frac{1}{\rho_{\rm g}}\nabla p+\frac{1}{\rho_{\rm g}}\nabla\cdot{\bf\it\tau}+{\bf\it g},$ (2)
with ${\bf\it g}=(0,-g)$, where $\rho_{\rm g}$, $\bf\it v$}, and $p$ denote the granular density, velocity, and pressure at position ${\bf\it r}=(X,Z)$ and time $t$, respectively.
In this study, we adopt the $\mu(I)$ rheology [14, 15] as the constitutive relation for the stress tensor $\bf\it\tau$}, which is expressed as
${\bf\it\tau}=\mu(I)\,p\,\frac{{\bf\it\dot{\gamma}}}{\dot{\gamma}},$ (3)
where the shear-rate tensor is defined by ${\bf\it\dot{\gamma}}=\nabla{\bf\it v}+(\nabla{\bf\it v})^{T}$. The effective friction coefficient $\mu(I)$ depends on the inertial number $I=\dot{\gamma}/\sqrt{p/\rho_{s}d^{2}}$, where $\dot{\gamma}$ denotes the second invariant of $\bf\it\dot{\gamma}$} and $\rho_{s}$ is the material density of the particles. Following Ref. 15, $\mu(I)$ is given by
$\mu(I)=\mu_{s}+\frac{\mu_{2}-\mu_{s}}{I_{0}/I+1},$ (4)
where $\mu_{s}$, $\mu_{2}$, and $I_{0}$ are material parameters reflecting the elasticity, dissipation, and friction of the particles. [14] At constant pressure $p$, this relation reduces to the Bingham-type model, which exhibits a finite yield stress below which no flow occurs. In this framework, granular materials can be regarded as non-Newtonian fluids whose viscosity depends on the local shear rate and pressure as
$\eta_{\mathrm{g}}(\dot{\gamma},p)=\frac{\mu(I)\,p}{\dot{\gamma}}.$ (5)
At the wall of the drum, the velocity satisfies
${\bf\it v}={\bf\it v}_{\mathrm{wall}},$ (6)
where
${\bf\it v}_{\mathrm{wall}}=-\Omega Z{\bf\it e}_{X}+\Omega X{\bf\it e}_{Z},$ (7)
and ${\bf\it e}_{X}$ and ${\bf\it e}_{Z}$ are the unit vectors along the $X$ and $Z$ axes, respectively. At the free surface of the granular bed, we impose the stress-free boundary conditions:
${\bf\it\tau}\cdot{\bf\it n}={\bf\it 0},$ (8)
where $\bf\it n$} denotes the unit normal vector to the free surface.
We employ the continuum model to theoretically derive the scaling law for the velocity profile, after confirming numerically that the model quantitatively reproduces the results of the DEM simulation. To compute the velocity field based on the continuum equations, we employ CFD with the finite-difference method, using a time step $\Delta t$ and grid width $\Delta x$, following Ref. 31, which studied an incompressible Newtonian fluid in a rotating container. Details of the CFD simulations are provided in Appendix C. The parameters $\mu_{s}$, $\mu_{2}$, and $I_{0}$ in Eq. (4) are estimated as $\mu_{s}=0.246$, $\mu_{2}=0.401$, and $I_{0}=0.133$ from the DEM simulation under uniform shear with the parameters given in Sec. II.1. The details of this estimation procedure are shown in Appendix D. Note that Eq. (4) is also used as the constitutive relation for grains interacting via the Hertzian contact model, with appropriate choices of these parameters. [14] The granular density is defined as $\rho_{\mathrm{g}}=\rho_{s}\phi$, where $\rho_{s}=m/(\pi d^{2}/4)$ is the material density. The packing fraction $\phi$ is set to $0.825$, corresponding to its value in the static regime of the rotating drum obtained from the DEM simulation (Appendix E). The spatial and temporal resolutions are set to $\Delta x=D/75$ and $\Delta t=1.0\times 10^{-4}\sqrt{d/g}$, respectively.
III Validation of the Continuum Model
To assess the validity of the continuum model introduced in Sec. II.2, we compare the velocity field in the steady state obtained from DEM simulations and from CFD calculations that numerically solve the continuum equations. Figures 2 (a) and (b) show the velocity field of granular material for $D=150d$ and $\Omega=6.28\times 10^{-3}\sqrt{g/d}$ obtained from the DEM and CFD simulations, respectively. The red solid lines in these figures represent the free surface. In both cases, the granular material convects in a counterclockwise direction. Near the wall, the flow exhibits the rigid rotation following the rotating drum. Near the free surface, there is a region where the velocity is almost parallel to the free surface.
To facilitate a quantitative comparison of the velocity field between DEM and CFD, we introduce a new Cartesian coordinate system $(x,z)$, where the $x$ axis is parallel to the free surface at $X=0$, as shown in Fig. 1. In this coordinate system, the depth from the free surface is represented by $z$. The coordinate transformation from $(X,Z)$ to $(x,z)$ is given by
@@ III Validation of the Continuum Model
where $\theta$ denotes the angle of the free surface at $X=0$. The position of the free surface in the DEM simulation is determined from the packing fraction, as detailed in Appendix E, and the free surface at $x=0$ is treated as $z=0$, which is numerically evaluated in Appendix F.
Next, we analyze the velocity $u$ in the $x$-direction as a function of $z$ at the center of the granular bed ($x=0$). Figures 3 show the velocity profiles for various values of $D$ and $\Omega$, obtained from DEM and CFD simulations. Open and closed symbols represent DEM and CFD results, respectively, which are quantitatively consistent. The velocity decreases with increasing $z$, approaching the rigid rotation profile $\Omega z$ at larger $z$, corresponding to the static flow regime, while significant deviations near $z=0$ correspond to the surface flow layer. The magnitude and slope of $u(z)$ in the surface flow layer increase with $D$ and $\Omega$, and the two regions are smoothly connected at the boundary where $u(z)\approx 0$. These results agree with Ref. 11 and demonstrate that the continuum model using the $\mu(I)$-rheology quantitatively reproduces the velocity field over a broad parameter space, beyond the limited conditions studied previously [20, 21, 22, 24].
Figure 2: Velocity fields obtained from the DEM (a) and CFD (b) simulations for $D=150d$ and $\Omega=1.26\times 10^{-3}\sqrt{g/d}$. The red solid line represents the free surface. The $x$-axis is taken parallel to the free surface, and the $z$-axis is normal to it.
Figure 3: (a) Velocity profile $u(z)$ at $x=0$ for $\Omega=6.28\times 10^{-3}\sqrt{g/d}$ with $D=150d$, $200d$, and $300d$. (b) Velocity profile $u(z)$ at $x=0$ for $D=150d$ with $\Omega=1.26\times 10^{-3}\sqrt{g/d}$ to $1.26\times 10^{-2}\sqrt{g/d}$. The open and closed symbols represent the results of the DEM and CFD simulations, respectively. The solid lines represent $\Omega z$ for the rigid rotation.
Next, to investigate the dependence of the velocity profile on $D$ and $\Omega$, we define the surface flow layer thickness $h$. Following Refs. 5, 8, we identify the point where $u(z)=0$ in Fig. 3 as the boundary between the surface flow layer and the underlying static bed, and estimate the surface flow layer thickness $h$ as the distance from the free surface to this boundary at $x=0$:
$u(z=h)=0.$ (10)
Figure 4 shows the surface flow layer thickness $h$ obtained from the DEM and CFD simulations as a function of $\Omega$ for different $D$. The numerical results from the DEM and CFD simulations are in good agreement. The surface flow layer thickness $h$ increases with $D$ and shows a weak dependence on $\Omega$. Although previous studies have estimated the boundary using different methods, [32, 33, 20, 11, 34] the choice of method does not affect the qualitative behavior of $h$, [5] which can be explained by the scaling of the velocity profile presented below (Appendix G).
IV Scaling laws
In this section, based on the continuum model introduced in Sec. II.2, we theoretically derive scaling laws for the velocity profile of granular flow in a rotating drum. We neglect the inertial term $\partial{\bf\it v}/\partial t$ in Eq. (2) by assuming steady flow. Using the drum diameter $D$, angular velocity $\Omega$, and granular density $\rho_{\mathrm{g}}$ as characteristic scales, we non-dimensionalize the variables as follows:
$\tilde{{\bf\it r}}=\frac{{\bf\it r}}{D},\;\;\tilde{\nabla}=D\nabla,\;\;\tilde{{\bf\it v}}=\frac{{\bf\it v}}{\Omega D},\;\;$ (11)
$\tilde{\dot{\gamma}}=\frac{\dot{\gamma}}{\Omega},\;\;\tilde{p}=\frac{p}{\rho_{\mathrm{g}}\Omega^{2}D^{2}},\;\;\tilde{\eta}_{\mathrm{g}}=\frac{\eta_{\mathrm{g}}}{\rho_{\mathrm{g}}\Omega D^{2}},$ (12)
where $\tilde{{\bf\it r}},\tilde{\nabla},\tilde{{\bf\it v}},\tilde{\dot{\gamma}},\tilde{p}$, and $\tilde{\eta}_{\mathrm{g}}$ are the dimensionless variables. We use $D$ as the characteristic length because we focus on the overall velocity profile in systems with $D\gg d$. Substituting these dimensionless variables into Eqs. (1)–(5), we obtain
$\tilde{\nabla}\cdot\tilde{{\bf\it v}}=0,$ (13)
$-\tilde{\nabla}\cdot(\tilde{{\bf\it v}}\tilde{{\bf\it v}})-\tilde{\nabla}\tilde{p}+\tilde{\nabla}\cdot\Big[\tilde{\eta}_{\mathrm{g}}\left(\tilde{\nabla}\tilde{{\bf\it v}}+(\tilde{\nabla}\tilde{{\bf\it v}})^{T}\right)\Big]-\frac{2}{Fr}{\bf\it e}_{Z}=0,$ (14)
$\tilde{\eta}_{\mathrm{g}}=\mu_{s}\frac{\tilde{p}}{\tilde{\dot{\gamma}}}+(\mu_{2}-\mu_{s})\frac{\tilde{p}}{I_{0}\sqrt{\phi\tilde{p}}/(d/D)+\tilde{\dot{\gamma}}}$, (15)
where the Froude number is defined as $Fr=\Omega^{2}D/(2g)$. We have summarized the dimensionless variables in Table 1. In the present system, the dimensionless parameters relevant for the velocity profile are $Fr$ and $d/D$. Thus, apart from the particle properties $\mu_{s}$, $\mu_{2}$, and $I_{0}$, the dimensionless velocity field $\tilde{{\bf\it v}}({\bf\it r})$ is determined solely by these two parameters. This scaling is consistent with the experiments reported in Ref. 35.
Table 1: Summary of dimensionless variables used in this study.
Dimensionless variables Definition / Physical meaning
$\tilde{\mathbf{r}}=\mathbf{r}/D$ Scaled position
$\tilde{\mathbf{v}}=\mathbf{v}/(\Omega D)$ Scaled velocity
$\tilde{p}=p/(\rho_{g}\Omega^{2}D^{2})$ Scaled pressure
$\tilde{\dot{\gamma}}=\dot{\gamma}/\Omega$ Scaled shear rate
$\tilde{\eta}_{g}=\eta_{g}/(\rho_{g}\Omega D^{2})$ Scaled granular viscosity
$Fr=\Omega^{2}D/(2g)$ Froude number
$d/D$ Particle-to-drum diameter ratio
$\mu_{s},\mu_{2},I_{0}$ Rheological parameters in the $\mu(I)$-rheology
Since in the previous section we have studied systems with $d/D\leq 1/150\ll 1$, we consider the limit $d/D\to 0$ in Eqs. (13)–(15). In this limit, the only remaining dimensionless parameter in these equations is the Froude number $Fr$. Hence, the solution of the continuum equations can be expressed as
where ${\bf\it V}\left(\tilde{{\bf\it r}};Fr\right)$ is a function that depends on $Fr$. Taking the limit $d/D\to 0$ in Eq. (15) is equivalent to treating the granular material as a simple plastic body with a yield stress proportional to the pressure, an approximation similar to that introduced in Ref. 36. From Eq. (16), the scaling law for the velocity profile $u(z)$ at $x=0$ is obtained as
$\frac{u(z)}{\Omega D}=U\left(\frac{z}{D};Fr\right),$ (17)
where $U\left(z/D;Fr\right)$ is a dimensionless scaling function.
Substituting Eq. (17) into the definition of the surface flow layer thickness, Eq. (10), we have
$U\left(\frac{h}{D};Fr\right)=0.$ (18)
This relation indicates that the dimensionless surface flow layer thickness $h/D$ is solely determined by the Froude number $Fr$, leading to the scaling law
$\frac{h}{D}=H\left(Fr\right),$ (19)
where $H(Fr)$ is a dimensionless scaling function. Thus, once the function $H(Fr)$ is known, the dependence of the surface flow layer thickness on $D$ and $\Omega$ is fully determined by the continuum model.
We verify the scaling laws using the numerical simulations. Figure 5 shows the normalized velocity profile $u/\Omega D$ as a function of $z/D$ for different $D/d$ values, with (a) $Fr=3.00\times 10^{-3}$ and (b) $Fr=1.20\times 10^{-2}$. Closed and open symbols represent the results of the CFD and DEM simulations, respectively, while the solid line corresponds to rigid rotation. Excellent collapse of the data is observed in Fig. 5, confirming the validity of the scaling law given in Eq. (17). Comparing the results for different $Fr$, we also find that the scaling function $U\left(z/D;Fr\right)$ exhibits a weak dependence on $Fr$.
Next, we plot $h/D$ against $Fr$ for various values of $d/D$ in Fig. 6. Closed and open symbols represent the results of the CFD and DEM simulations, respectively. This figure confirms the validity of the scaling law, Eq. (19). We have focused on the steady flow regime with low $Fr<10^{-2}$. In this regime, the scaling function $H(Fr)$ in Eq. (19) can be expanded as
$H\left(Fr\right)=H^{(0)}+FrH^{(1)}+O(Fr^{2}),$ (20)
where $H^{(0)}$ and $H^{(1)}$ are constants. As shown in Appendix H, $H^{(0)}>0$ is required from flux conservation in the steady flow. The solid line in Fig. 6 represents Eq. (20) with $H^{(0)}=1.1$ and $H^{(1)}=1.2$, estimated via the least squares method. The agreement of the data with the solid line further supports the validity of the scaling law. The scaling laws (19) and (20) derived from the continuum model indicate that $h/D$ is nearly constant for $Fr\ll 1$, as shown in Fig. 6. Accordingly, the surface flow layer thickness $h$ is proportional to $D$ but independent of $\Omega$ in this regime. This dependence is consistent with previous experimental results. [2]
Figure 6: Normalized surface flow layer thickness $h/D$ against the Froude number FrFr for various values of $D$. The open and closed symbols represent the results of the DEM and CFD simulations, respectively. The solid line represents Eq. (20) with $H^{(0)}=1.1$ and $H^{(1)}=1.2$, which are estimated by the least squares method.
V Dicussion and Conclusion
In this paper, we aimed to understand how the surface flow layer thickness $h$ and the velocity profile of granular material in a two-dimensional rotating drum depend on the drum diameter $D$ and angular velocity $\Omega$. To achieve this, we introduced a continuum model based on the $\mu(I)$-rheology and performed numerical simulations (CFD), complemented by discrete element method (DEM) simulations for validation. This combined approach allowed us to quantitatively reproduce the velocity fields observed in the DEM simulations and to capture the overall flow behavior, including the rigid rotation in the static regime near the drum wall and the flow parallel to the free surface in the surface layer. From dimensional analysis of the continuum equations, we analytically derived scaling laws for the velocity profile (Eq. (17)) and the surface layer thickness (Eq. (19)). The numerical results confirm the validity of these scaling laws, demonstrating that the surface flow thickness $h$ is proportional to $D$ and nearly independent of $\Omega$ for sufficiently small Froude numbers. Therefore, we conclude that the derived scaling laws are valid for 2D flows in half-filled rotating drums with large $D/d$ and low Froude numbers ($10^{-4}<Fr<10^{-2}$), where the local and incompressible $\mu(I)$-rheology provides an appropriate continuum description.
The $\mu(I)$-rheology has been shown to be applicable to both two-dimensional [14] and three-dimensional granular systems [15, 16]. When applying a continuum model based on the $\mu(I)$-rheology to granular flows in three-dimensional rotating drums, the scaling laws derived in this study, Eqs. (17) and (19), are expected to hold provided that the flow is approximately uniform along the axial direction. In such cases, the same dimensional arguments apply, although the scaling functions $U(z/D;Fr)$ and $H(Fr)$ may differ quantitatively due to changes in packing fraction and rheological parameters between 2D and 3D systems.
The scaling law for the surface flow layer thickness, Eq. (19), predicts $h\sim\mathrm{const.}$ in the limit $\Omega\to 0$, which is consistent with the experimental results reported in Ref. 2, although the estimated value of $H^{(0)}$ differs from that obtained in our two-dimensional simulations. By contrast, several experimental studies using axially short rotating drums [3, 5, 6, 8] have reported a power-law dependence $h\sim\Omega^{\alpha}$, based on phenomenological assumptions such as a constant shear rate.
Our preliminary three-dimensional simulations indicate that, for sufficiently long rotating cylinders, the flow remains nearly uniform along the axial direction except in narrow regions near the sidewalls. In this regime, the scaling laws derived from the two-dimensional analysis are effectively satisfied. For shorter cylinders, however, sidewall friction induces pronounced axial inhomogeneities, leading to deviations from the predicted scaling behavior. These results suggest that sidewall effects [37] play a key role in the breakdown of the scaling laws and may account for the discrepancies between our predictions and experiments conducted in axially thin drums. Further investigation is required to quantitatively characterize these effects and to extend the theoretical framework to fully three-dimensional geometries.
In this study, we considered a rotating drum half-filled with granular material. Nevertheless, similar flow patterns as shown in Fig. 1 are generally observed for a range of fill fractions in previous studies. We have not explicitly investigated the dependence of the velocity field on the fill fraction, and the derivation and confirmation of scaling laws for different fill fractions will be left for future work.
We have used the local $\mu(I)$-rheology with incompressible conditions in the continuum model. Recent studies show that continuum models with the local, incompressible $\mu(I)$-rheology may exhibit ill-posed behavior, particularly in the low- and high-inertial-number ($I$) regimes [38, 18]. In our simulations, the high-$I$ regime does not occur, and for the low-$I$ region, a numerical cutoff is introduced for the effective viscosity $\eta_{\mathrm{g}}$ (see Appendix C.2), which mitigates the ill-posedness. Nevertheless, further investigation is needed to fully clarify the dependence on numerical parameters.
The $\mu(I)$-rheology has been extended to include non-local effects, which lead to deviations from a simple local constitutive law and generate exponential creep tails near the interface between the surface flow layer and the static flow regime [17, 18, 19]. In the present study, such effects are neglected. Analytical and numerical studies have shown that non-local effects become negligible for sufficiently large systems in granular flows between rough parallel plates [39]. Although an analogous analytical demonstration is not yet available for rotating drums, we expect that a similar argument applies in the limit of large drums with sufficiently large $D/d$. Consistent with this expectation, the scaling laws derived from the local $\mu(I)$-rheology are confirmed in our simulations for two-dimensional rotating drums with $D/d\geq 150$. In contrast, for smaller drums or three-dimensional systems with frictional side walls, non-local effects are expected to play a more significant role and may modify the scaling laws. Such effects could contribute to the discrepancies between our predictions and the experimental results reported in Refs. 3, 5, 6, 8. Further studies are required to clarify the influence of non-local effects and sidewall friction on the velocity field in rotating drums.
Although we have obtained the scaling laws through non-dimensionalization, the explicit forms of the scaling functions have not yet been derived analytically. Analytical studies on related systems provide useful guidance: for example, Haji-Sheikh and Elliot [40] calculated the velocity field of an incompressible Newtonian fluid in a rotating drum, and Gray et al. [41] derived the streamlines of granular flow using a simplified constitutive law. Extending such analytical approaches to the present system could offer a deeper theoretical understanding of granular flows and help to explicitly determine the scaling functions.
In conclusion, the novelty of the present work lies not merely in reproducing velocity fields observed in experiments or DEM simulations, but in providing a theoretical framework that clarifies how the surface flow layer thickness $h$ and the velocity profile in rotating drums depend on the system parameters. By applying dimensional analysis to the continuum equations based on the local, incompressible $\mu(I)$-rheology, we demonstrate that, in the limit of large $D/d$ and sufficiently small Froude numbers, the two-dimensional flow is governed primarily by the Froude number, leading to scaling laws in which the surface layer thickness is proportional to the drum diameter and nearly independent of the angular velocity. This result helps to rationalize apparent discrepancies among previous experimental and numerical studies by suggesting that reported power-law dependencies on $\Omega$ arise from additional dimensionless parameters not captured in the idealized two-dimensional limit, such as sidewall friction, finite axial length, or non-local rheological effects. In this sense, the present scaling analysis provides new physical insight into the origin and limitations of commonly observed scaling behaviors in granular flows within rotating drums and offers a systematic basis for interpreting deviations from the ideal scaling regime.
Acknowledgements.
The authors thank S. Goto, Y. Motoori, D. Watanabe, and I. Procaccia for fruitful discussions. We would like to thank Editage (www.editage.jp) for English language editing. This study was supported by JSPS KAKENHI under Grant No. JP23K03248 and JP21H01006.
Author declarations
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Hiroki Oba: Conceptualization (equal); Data curation (lead); Formal Analysis (lead); Methodology (equal); Software (lead); visualization (lead); Writing – original draft (lead). Michio Otsuki Conceptualization (equal); Methodology (equal); Supervision (lead); Writing – review & editing (lead).
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Details of DEM simulations